# DualCodec: A Low-Frame-Rate, Semantically-Enhanced Neural Audio Codec for Speech Generation

### v1

@@ Abstract
Neural audio codecs form the foundational building blocks for language model (LM)-based speech generation. Typically, there is a trade-off between frame rate and audio quality. This study introduces a low-frame-rate, semantically enhanced codec model. Existing approaches distill semantically rich self-supervised (SSL) representations into the first-layer codec tokens. This work proposes DualCodec, a dual-stream encoding approach that integrates SSL and waveform representations within an end-to-end codec framework. In this setting, DualCodec enhances the semantic information in the first-layer codec and enables the codec system to maintain high audio quality while operating at a low frame rate. Note that a low-frame-rate codec improves the efficiency of speech generation. Experimental results on audio codec and speech generation tasks confirm the effectiveness of the proposed DualCodec compared to state-of-the-art codec systems, such as Mimi Codec, SpeechTokenizer, DAC, and Encodec. Demos and codes are available at: https://dualcodec.github.io.
@@ 1 Introduction
  • Audio quality: A high codec reconstruction quality is essential to TTS’s audio quality [10]. This becomes challenging for low-frame-rate audio codecs operating at a low bitrate.
Figure 1: The dual encoding method for neural audio codecs. The upper stream is SSL encoding, and the lower stream is waveform encoding. Given a speech input, the SSL feature is obtained from a pretrained w2v-bert-2 [14] model and then encoded as the codec’s first-layer token (RVQ-1). The remaining RVQ layers encodes the residual between the waveform feature and the RVQ-1 feature, and outputs audio. The framework is trained end-to-end requiring an additional L2 SSL feature loss in addition to codec training losses.
@@
We argue that our three summarized design principles can be fulfilled in a single framework: incorporating an explicit semantic-related codec layer, maintaining a low frame rate, and preserving high audio quality. To achieve this, we propose DualCodec, which unifies SSL and waveform representations in a dual encoding framework. In this framework, the tokens from the first layer are semantically enhanced directly from SSL feature. DualCodec achieves a low frame rate while maintaining exceptional audio quality, with benefits from dual encoding, low-frame-rate configuration, larger codebook sizes, and DAC [12] codec innovations. We will open-source the DualCodec in both 12.5Hz and 25Hz settings, as well as the DualCodec-based TTS systems in this work. To the best of our knowledge, this is the first open-source 12.5 Hz low frame-rate codec [^1] The Mimi codec is the first open-weight 12.5Hz codec, but its data and training codes are not available. We train on a public dataset, and has released our model weights and training code., and the first open-source 12.5Hz TTS.
@@ 2.2 Neural Audio Codecs
Semantic Enhancement.  Enhancing the semantic information in audio codec tokens can benefit SLM intelligibility and stability, especially in the RVQ-1 token where the AR generation can accumulate prediction errors [9]. SpeechTokenizer [9] proposed to enhance RVQ-1 through semantic distillation. The approach introduces a distillation loss between the RVQ-1 codebook vector and SSL feature extracted from the same speech input. Mimi codec [11] applied the semantic distillation on a 12.5Hz codec. Concurrent work X-Codec [21] concatenates SSL representations with waveform features before RVQ, and jointly optimizes acoustic and semantic reconstruction objectives [22]. Our dual encoding method differs by encoding RVQ-1 and RVQ-rest tokens from different sources, closely aligning with SLM’s AR+NAR paradigm. Meanwhile, our work explores low frame rate, audio quality improvement, and multilingual evaluation that are not explored in the concurrent work.
Low Bitrate and Low Frame Rate.  Vanilla neural audio codecs operate at more than 4kbps bitrate and above 50Hz frame rates [^2] Frame rate is the number of segments per second in the codec’s latent representation. Bitrate indicates the data amount per unit of time. [2, 3]. Some recent efforts focus on designing low bitrate codec systems, benefiting the TTS efficiency by reducing the token length. WavTokenizer [23], SingleCodec [24] and TS3Codec [1] used a single VQ layer with a high frame rate. Mimi codec [11] reduces frame rate to 12.5Hz while keeping multiple RVQ layers. The latter low-frame-rate approach has better TTS efficiency by effectively shortening the sequence length in AR part. However, we still find the previous low-frame-rate codec has unsatisfiable audio quality especially at low bitrates $<$1kbps.
@@ 3 Method
The two streams are jointly optimized. By using the two encoding streams, we obtain semantic-rich RVQ-1 tokens, with remaining layers (RVQ-rest) focused on the remaining acoustic aspects in the waveform feature. This “disentanglement” is achieved by subtracting RVQ-1 feature from waveform feature, before obtaining RVQ-rest tokens. Finally to decode audio, the RVQ-1 feature is re-summed to the codebook vectors of RVQ-rest. During TTS training, both encoding streams are required to obtain training tokens. During TTS inference, only the codec decoder is used to produce audio.
@@ 3.1 SSL Encoding
Downsampler.  We downsample the 50Hz SSL feature into our codec’s frame rate using simple 1D average pooling, using kenel_size = stride_size = downsampling_factor. The downsampling_factor is 2 for 25Hz target frame rate, and is 4 for 12.5Hz.
ResNet Encoder and Decoder.  These networks are used to process the SSL feature before and after the VQ module. This allows the VQ tokens to capture more complex semantic patterns. The decoder mirrors the encoder. Both contain stacked ConvNeXt [29], which are the latest ResNet [30] variants. Each network has 13M parameters. There’s no down-sampling or up-sampling operation in these ResNet modules.
VQ Module.  This module discretizes the ResNet Encoder output $\mathbf{Z}_{\text{ssl}}\in\mathbb{R}^{H\times T}$ into a 1D token sequence $RVQ\_1\in\mathbb{Z}^{1\times T}$, where H is the hidden dimension and T is the feature length. Following [12], $RVQ\_1$ is computed by finding the closest codebook vector to the projected input: $RVQ\_1=\arg\min_{k}||\ell_{2}(W_{\text{in}}\mathbf{Z}_{ssl})-\ell_{2}(e_{k})||_{2}$. Here, $W_{\text{in}}\in\mathbb{R}^{D\times H}$ is the input projection matrix with $D=8$,$H=1024$, $\ell_{2}$ is the L2-normalizaton, $e_{1},e_{2},...,e_{k}$ are codebook vectors, $e_{k}\in\mathbb{R}^{H\times T}$. We obtain RVQ-1-feature after processing with the ResNet decoder: $RVQ\_1\_feat=\text{ResNet}(e_{k})$.
@@ 3.2 Waveform Encoding
RVQ module.  This module has $N-1$ layers of VQ. Each VQ layer quantizes the residual error of the previous layer [3]. The input to this module is the residual between the waveform feature and the $RVQ\_1\_feat$. It discretizes into RVQ_rest $\in\mathbb{Z}^{(N-1)\times T}$. After obtaining the RVQ_rest tokens, each selected codebook vector $e_{k}$ is added together with $RVQ\_1\_feat$. This continuous feature summarizes SSL encoding and waveform encoding, and is used as input to the codec decoder. Notably, we employ RVQ dropout [2] during training. That is, we only use the first $q$ RVQ quantizers each time, where $q\in[0,N-1]$ is randomly chosen. When $q=0$, only the SSL encoding stream is used, and the model only decodes the RVQ-1 tokens.
@@ 4.1 Model training setup
We use the 100K-hour multilingual, 24kHz speech dataset Emilia [32] in Amphion toolkit [33], and 8 A100 GPUs for training. Each codec model is trained for 1 epoch (400K steps) with a 3-second segment length. Each TTS model is trained for 2 epochs (around 600K steps). There are $N=8$ codebook layers in DualCodec. In the following sections, we comprehensively evaluate our codec’s semantic content WER, audio reconstruction quality, TTS performance and speed, respectively.
@@ 4.2 Semantic content analysis
Config WER(%$\downarrow$ WER(%$\downarrow$
@@
Group A baselines.  Group A (A1 and A2) models are ground-truth and official SpeechTokenizer checkpoint, respectively. The A2 SpeechTokenizer model has semantic distillation and 1024 EMA codebooks. While it is trained on English-only data, we report its Chinese performance and find that its RVQ-1 has extremely high Chinese WER. Listening to samples, we suggest that the model’s RVQ-1 tokens, although having semantic distillation, lacks pitch information. Pitch is critical for understanding Chinese.
@@ 4.3 Audio quality analysis
DualCodec.  Group G highlights DualCodec’s performance under ablated settings.
1. Effect of Dual Encoding.  Models labeled G1 and G3 have same RVQ codebook settings as Group F models, allowing us to ablate on the impact of dual encoding. Comparing G1 (DualCodec 25Hz) with F1 (DAC-repro 25Hz), G1 achieves similar performance across most objective metrics but demonstrates a significant improvement in UTMOS (3.86 vs. 3.29) and MUSHRA (79.5 vs. 68.8), indicating DualCodec’s noticeable enhancement in perceptual audio quality. Comparing G3 (DualCodec 12.5Hz) with F2 (DAC-repro 12.5Hz) shows a similar trend where DualCodec has higher perceptual metrics UTMOS (3.94 vs. 3.87) and MUSHRA (83.5 vs. 81.8). We conclude that DualCodec’s additional semantic encoding stream can enhance the perceptual audio quality.
@@ 4.4 TTS analysis
Performance Comparisons.  Table 4 demonstrates that DualCodec outperforms SpeechTokenizer and Mimi baselines in both TTS. We attribute the superior WER performance to DualCodec’s more accurate semantic content, and the SIM to DualCodec’s better codec reconstruction quality. The AR+SoundStorm TTS paired with DualCodec 25Hz achieves the best performance, followed by DualCodec 12.5Hz. When comparing RTF, the trade-off between quality and inference speed becomes apparent: DualCodec 25Hz consistently achieves lower WER and higher SIM scores, making it the ideal choice for tasks prioritizing accuracy and similarity. On the other hand, DualCodec 12.5Hz provides much faster inference at the cost of acceptable degrees of performance decrease. We also notice that Mimi and SpeechTokenizer have excessively large Chinese WERs in AR+SoundStorm. We suggest this is due to a lack of RVQ-1 semantic pitch information, which becomes more notable in AR+SoundStorm TTS because its NAR does not have text prompting for extracting the necessary pitch.
@@ 5 Conclusion
We introduced DualCodec, a low-frame-rate, semantically-enhanced neural audio codec designed for efficient speech generation. By leveraging dual encoding, low frame rates and larger codebooks, DualCodec improves semantic accuracy, audio reconstruction quality, TTS performance and efficiency. A limitation in this work is the gap between the 12.5Hz DualCodec-based TTS and the 25Hz ones. Future works could explore methods to further increase the 12.5Hz codec’s semantic accuracy and quality upper-bound. DualCodec also has the potential to be used in real-time multimodal LLMs, since it allows both efficient audio encoding and high quality audio reconstruction.
@@ 6 Acknowledgements
This projected is funded by 2023 Shenzhen stability Science Program.
@@ Appendix B The Inference Speed of DualCodec
Results.  From Table 5, we observe that DualCodec’s RTFs, while not the most competitive among the evaluated systems, remain very acceptable given its design and capabilities. For instance, DualCodec 25Hz achieves an encoding RTF of 0.015 and a decoding RTF of 0.002, and DualCodec 12Hz attains an encoding RTF of 0.015 and a decoding RTF of 0.003. Importantly, the decoding phase of DualCodec remains highly efficient, with RTF values on par with or better than most baselines. This ensures the DualCodec-based TTS systems remains lightweight and fast during deployment.
@@ Appendix C More Open-Source DualCodec TTS Systems
Table 6: A summary of speech language model performance on Seed-TTS-Eval. We add “DualCodec-” prefix for DualCodec-based models. We are open-sourcing DualCodec-VALLE, DualCodec-MaskGCT, and DualCodec-FlattenedAR. See the demo page link for more details.
@@
We are open sourcing more TTS systems based on DualCodec, trained on the 100K-hour Emilia [32] dataset. Apart from the VALL-E and AR+SoundStorm TTS, we also demonstrate DualCodec’s effectiveness for an AR-only FlattenAR architecture, and the MaskGCT [8] architecture.
@@
We plan to continue developing more DualCodec-based TTS.

### v2

@@ Abstract
Neural audio codecs form the foundational building blocks for language model (LM)-based speech generation. Typically, there is a trade-off between frame rate and audio quality. This study introduces a low-frame-rate, semantically enhanced codec model. Existing approaches distill semantically rich self-supervised (SSL) representations into the first-layer codec tokens. This work proposes DualCodec, a dual-stream encoding approach that integrates SSL and waveform representations within an end-to-end codec framework. In this setting, DualCodec enhances the semantic information in the first-layer codec and enables the codec system to maintain high audio quality while operating at a low frame rate. Note that a low-frame-rate codec improves the efficiency of speech generation. Experimental results on audio codec and speech generation tasks confirm the effectiveness of the proposed DualCodec compared to state-of-the-art codec systems, such as Mimi Codec, SpeechTokenizer, DAC, and Encodec. Demos are available at: https://dualcodec.github.io, code is available at: https://github.com/jiaqili3/DualCodec
@@ 1 Introduction
  • Audio quality: A high codec reconstruction quality is essential to TTS's audio quality [10]. This becomes challenging for low-frame-rate audio codecs operating at a low bitrate.
Figure 1: The dual encoding method for neural audio codecs. The upper stream is SSL encoding, and the lower stream is waveform encoding. Given a speech input, the SSL feature is obtained from a pretrained w2v-bert-2 [14] model and then encoded as the codec's first-layer token (RVQ-1). The remaining RVQ layers encodes the residual between the waveform feature and the RVQ-1 feature, and outputs audio. The framework is trained end-to-end requiring an additional L2 SSL feature loss in addition to codec training losses.
@@
We argue that our three summarized design principles can be fulfilled in a single framework: incorporating an explicit semantic-related codec layer, maintaining a low frame rate, and preserving high audio quality. To achieve this, we propose DualCodec, which unifies SSL and waveform representations in a dual encoding framework. In this framework, the tokens from the first layer are semantically enhanced directly from SSL feature. DualCodec achieves a low frame rate while maintaining exceptional audio quality, with benefits from dual encoding, low-frame-rate configuration, larger codebook sizes, and DAC [12] codec innovations. We will open-source the DualCodec in both 12.5Hz and 25Hz settings, as well as the DualCodec-based TTS systems in this work. To the best of our knowledge, this is the first open-source 12.5 Hz low frame-rate codec [^1] The Mimi codec is the first open-weight 12.5Hz codec, but its data and training code are not available. We train on a public dataset, and have released our model weights and training code (see abstract)., and the first open-source 12.5Hz TTS.
@@ 2.2 Neural Audio Codecs
Semantic Enhancement.  Enhancing the semantic information in audio codec tokens can benefit SLM intelligibility and stability, especially in the RVQ-1 token where the AR generation can accumulate prediction errors [9]. SpeechTokenizer [9] proposed to enhance RVQ-1 through semantic distillation. The approach introduces a distillation loss between the RVQ-1 codebook vector and SSL feature extracted from the same speech input. Mimi codec [11] applied the semantic distillation on a 12.5Hz codec. Concurrent work X-Codec [21] concatenates SSL representations with waveform features before RVQ, and jointly optimizes acoustic and semantic reconstruction objectives [22]. Our dual encoding method differs by encoding RVQ-1 and RVQ-rest tokens from different sources, closely aligning with SLM's AR+NAR paradigm. Meanwhile, our work explores low frame rate, audio quality improvement, and multilingual evaluation that are not explored in the concurrent work.
Low Bitrate and Low Frame Rate.  Vanilla neural audio codecs operate at more than 4kbps bitrate and above 50Hz frame rates [^2] Frame rate is the number of segments per second in the codec’s latent representation. Bitrate indicates the data amount per unit of time. [2, 3]. Some recent efforts focus on designing low bitrate codec systems, benefiting the TTS efficiency by reducing the token length. WavTokenizer [23], SingleCodec [24] and TS3Codec [1] used a single VQ layer with a high frame rate. Mimi codec [11] reduces frame rate to 12.5Hz while keeping multiple RVQ layers. The latter low-frame-rate approach has better TTS efficiency by effectively shortening the sequence length in AR part. However, we still find the previous low-frame-rate codec has unsatisfiable audio quality especially at low bitrates <<1kbps.
@@ 3 Method
The two streams are jointly optimized. By using the two encoding streams, we obtain semantic-rich RVQ-1 tokens, with remaining layers (RVQ-rest) focused on the remaining acoustic aspects in the waveform feature. This ``disentanglement" is achieved by subtracting RVQ-1 feature from waveform feature, before obtaining RVQ-rest tokens. Finally to decode audio, the RVQ-1 feature is re-summed to the codebook vectors of RVQ-rest. During TTS training, both encoding streams are required to obtain training tokens. During TTS inference, only the codec decoder is used to produce audio.
@@ 3.1 SSL Encoding
Downsampler.  We downsample the 50Hz SSL feature into our codec's frame rate using simple 1D average pooling, using kenel_size = stride_size = downsampling_factor. The downsampling_factor is 2 for 25Hz target frame rate, and is 4 for 12.5Hz.
ResNet Encoder and Decoder.  These networks are used to process the SSL feature before and after the VQ module. This allows the VQ tokens to capture more complex semantic patterns. The decoder mirrors the encoder. Both contain stacked ConvNeXt [29], which are the latest ResNet [30] variants. Each network has 13M parameters. There's no down-sampling or up-sampling operation in these ResNet modules.
[…]
@@ 3.2 Waveform Encoding
RVQ module.  This module has N1N-1 layers of VQ. Each VQ layer quantizes the residual error of the previous layer [3]. The input to this module is the residual between the waveform feature and the RVQ_1_featRVQ\_1\_feat. It discretizes into RVQ_rest $\in\mathbb{Z}^{(N-1)\times T}$. After obtaining the RVQ_rest tokens, each selected codebook vector $e_{k}$ is added together with RVQ_1_featRVQ\_1\_feat. This continuous feature summarizes SSL encoding and waveform encoding, and is used as input to the codec decoder. Notably, we employ RVQ dropout [2] during training. That is, we only use the first $q$ RVQ quantizers each time, where $q\in[0,N-1]$ is randomly chosen. When $q=0$, only the SSL encoding stream is used, and the model only decodes the RVQ-1 tokens.
@@ 4.1 Model training setup
We use the 100K-hour multilingual, 24kHz speech dataset Emilia [32] in Amphion toolkit [33], and 8 A100 GPUs for training. Each codec model is trained for 1 epoch (400K steps) with a 3-second segment length. Each TTS model is trained for 2 epochs (around 600K steps). There are $N=8$ codebook layers in DualCodec. In the following sections, we comprehensively evaluate our codec's semantic content WER, audio reconstruction quality, TTS performance and speed, respectively.
@@ 4.2 Semantic content analysis
Config $\downarrow WER()\downarrow$
@@
Group A baselines.  Group A (A1 and A2) models are ground-truth and official SpeechTokenizer checkpoint, respectively. The A2 SpeechTokenizer model has semantic distillation and 1024 EMA codebooks. While it is trained on English-only data, we report its Chinese performance and find that its RVQ-1 has extremely high Chinese WER. Listening to samples, we suggest that the model's RVQ-1 tokens, although having semantic distillation, lacks pitch information. Pitch is critical for understanding Chinese.
@@ 4.3 Audio quality analysis
DualCodec.  Group G highlights DualCodec's performance under ablated settings.
1. Effect of Dual Encoding.  Models labeled G1 and G3 have same RVQ codebook settings as Group F models, allowing us to ablate on the impact of dual encoding. Comparing G1 (DualCodec 25Hz) with F1 (DAC-repro 25Hz), G1 achieves similar performance across most objective metrics but demonstrates a significant improvement in UTMOS (3.86 vs. 3.29) and MUSHRA (79.5 vs. 68.8), indicating DualCodec's noticeable enhancement in perceptual audio quality. Comparing G3 (DualCodec 12.5Hz) with F2 (DAC-repro 12.5Hz) shows a similar trend where DualCodec has higher perceptual metrics UTMOS (3.94 vs. 3.87) and MUSHRA (83.5 vs. 81.8). We conclude that DualCodec’s additional semantic encoding stream can enhance the perceptual audio quality.
@@ 4.4 TTS analysis
Performance Comparisons.  Table 4 demonstrates that DualCodec outperforms SpeechTokenizer and Mimi baselines in both TTS. We attribute the superior WER performance to DualCodec’s more accurate semantic content, and the SIM to DualCodec's better codec reconstruction quality. The AR+SoundStorm TTS paired with DualCodec 25Hz achieves the best performance, followed by DualCodec 12.5Hz. When comparing RTF, the trade-off between quality and inference speed becomes apparent: DualCodec 25Hz consistently achieves lower WER and higher SIM scores, making it the ideal choice for tasks prioritizing accuracy and similarity. On the other hand, DualCodec 12.5Hz provides much faster inference at the cost of acceptable degrees of performance decrease. We also notice that Mimi and SpeechTokenizer have excessively large Chinese WERs in AR+SoundStorm. We suggest this is due to a lack of RVQ-1 semantic pitch information, which becomes more notable in AR+SoundStorm TTS because its NAR does not have text prompting for extracting the necessary pitch.
@@ 5 Conclusion
We introduced DualCodec, a low-frame-rate, semantically-enhanced neural audio codec designed for efficient speech generation. By leveraging dual encoding, low frame rates and larger codebooks, DualCodec improves semantic accuracy, audio reconstruction quality, TTS performance and efficiency. A limitation in this work is the gap between the 12.5Hz DualCodec-based TTS and the 25Hz ones. Future works could explore methods to further increase the 12.5Hz codec's semantic accuracy and quality upper-bound. DualCodec also has the potential to be used in real-time multimodal LLMs, since it allows both efficient audio encoding and high quality audio reconstruction.
@@ 6 Acknowledgements
The work is partially supported by the NSFC under Grant 62376237; the Shenzhen Science and Technology Program ZDSYS20230626091302006; 2023 Shenzhen Stability Science Program; Program for Guangdong Introducing Innovative and Entrepreneurial Teams 2023ZT10X044.
@@ Appendix B The Inference Speed of DualCodec
Results.  From Table 5, we observe that DualCodec's RTFs, while not the most competitive among the evaluated systems, remain very acceptable given its design and capabilities. For instance, DualCodec 25Hz achieves an encoding RTF of 0.015 and a decoding RTF of 0.002, and DualCodec 12Hz attains an encoding RTF of 0.015 and a decoding RTF of 0.003. Importantly, the decoding phase of DualCodec remains highly efficient, with RTF values on par with or better than most baselines. This ensures the DualCodec-based TTS systems remains lightweight and fast during deployment.
@@ Appendix C More Open-Source DualCodec TTS Systems
Table 6: A summary of speech language model performance on Seed-TTS-Eval. We add ``DualCodec-'' prefix for DualCodec-based models. We are open-sourcing DualCodec-VALLE, DualCodec-MaskGCT, and DualCodec-FlattenedAR. See the demo page link for more details.
@@
We are open sourcing more TTS systems based on DualCodec, trained on the 100K-hour Emilia [32] dataset. Apart from the VALL-E and AR+SoundStorm TTS, we also demonstrate DualCodec's effectiveness for an AR-only FlattenAR architecture, and the MaskGCT [8] architecture.